OpenStudy (anonymous):

Y over -2 +5=-4 solve y

OpenStudy (anonymous):

subtract 5 to get (y/-2) = -9. multiply by -2 to get y = 18

OpenStudy (lgbasallote):

\[\frac{y}{-2} + 5 = -4\] \[\frac{y}{-2} = -4 - 5\] \[\frac{y}{-2} = -9\] \[y = -9 \times -2\] \[y = 18\]